# On the Analysis of Misspecified Variational Inequalities with Nonlinear Constraints

### v1

@@ 5 Numerical Experiment
Using the above inequality in (C), multiplying both sides by $2\rho$ and rearranging the terms, we obtain
$\left(1-\rho L^{\Phi}_{\lambda\theta}\alpha_{k}\right)\|\lambda_{k+1}-\lambda^{*}\|^{2}\leq\|\lambda_{k}-\lambda^{*}\|^{2}+2\rho\left(T_{k}+R_{k}\right),$ (27)
$\text{ where }T_{k}$ $\triangleq\tfrac{1}{2\gamma}\left(\|x_{k}-x^{*}\|^{2}-\|x_{k+1}-x^{*}\|^{2}\right)+\tfrac{L_{F,x}}{2}\left(\|x_{k}-x_{k-1}\|^{2}-\|x_{k+1}-x_{k}\|^{2}\right)$
$\quad+\langle r_{k+1},x_{k+1}-x^{*}\rangle-\langle r_{k},x_{k}-x^{*}\rangle,$
$R_{k}$ $\triangleq L_{F,\theta}\|\theta_{k+1}-\theta^{*}\|\|x_{k+1}-x^{*}\|+L^{\Phi}_{x\theta}\|\theta_{k}-\theta^{*}\|\big(\|x_{k+1}-x_{k}\|+\|x_{k}-x^{*}\|\big)$
$\quad+\tfrac{L_{F,\theta}}{2}\|\theta_{k}-\theta_{k-1}\|^{2}+\tfrac{L^{\Phi}_{\lambda\theta}}{2\alpha_{k}}\|\theta_{k}-\theta^{*}\|^{2}+\rho L_{f,\theta}\|\theta_{k}-\theta^{*}\|\sum_{j\in J_{k}^{+}}|f_{j}(x_{k+1},\theta^{*})|.$
Since $0<\alpha_{k}<1$ and $0<\rho L_{\lambda\theta}^{\Phi}\leq 1$, dividing both sides of (27) by $\left(1-\rho L^{\Phi}_{\lambda\theta}\alpha_{k}\right)$ we get
$\|\lambda_{k+1}-\lambda^{*}\|^{2}\leq\frac{1}{\left(1-\rho L^{\Phi}_{\lambda\theta}\alpha_{k}\right)}\|\lambda_{k}-\lambda^{*}\|^{2}+\frac{2\rho}{\left(1-\rho L^{\Phi}_{\lambda\theta}\alpha_{k}\right)}\left(T_{k}+R_{k}\right).$ (28)
Summing $T_{k}$ over $k=0,\dots,K-1$, we get
@@
From the definition of $T_{k}$ and telescoping with $x_{-1}=x_{0}$ and $r_{0}=0$ in (C), we have
@@
Choose $\gamma\in\Big(0,\min\{(L^{\Phi}+2L_{F,x}+L_{F,\theta})^{-1},(L_{F,x}+L_{F,\theta})^{-1}\}\Big)$. With this choice, the coefficient of $\|x_{K}-x^{*}\|^{2}$ in the bound (30) for $\sum_{k=0}^{K-1}T_{k}$ is non-positive and may be discarded. Hence $\sum_{k=0}^{K-1}T_{k}\leq\frac{1}{2\gamma}\|x_{0}-x^{*}\|^{2}+\frac{L_{F,\theta}}{2}\|\theta_{K}-\theta_{K-1}\|^{2}.$ By compactness of $\Theta$, $\|\theta_{K}-\theta_{K-1}\|\leq\|\theta_{K}-\theta^{*}\|+\|\theta_{K-1}-\theta^{*}\|\leq 2D_{\Theta}$, hence
@@
where $C_{T}<\infty$ is independent of $K$. In particular, $\sum_{k=0}^{\infty}T_{k}\leq C_{T}<\infty$.
Again, by compactness, $\|x_{k+1}-x^{*}\|\leq D_{x}$ and $\|x_{k+1}-x_{k}\|\leq\|x_{k+1}-x^{*}\|+\|x_{k}-x^{*}\|\leq 2D_{x}$. Continuity on compact $X$ gives $\sum_{j=1}^{J}|f_{j}(x,\theta^{*})|\leq JD_{f}(\theta^{*})$ for all $x\in X$, hence $\sum_{j\in J_{k}^{+}}|f_{j}(x_{k+1},\theta^{*})|\leq JD_{f}(\theta^{*})$. Therefore, for every $k$
@@
$\quad+\frac{L_{\lambda\theta}^{\Phi}}{2\alpha_{k}}\|\theta_{k}-\theta^{*}\|^{2}+\rho L_{f,\theta}JD_{f}(\theta^{*})\|\theta_{k}-\theta^{*}\|$
$\leq L_{F,\theta}D_{x}\|\theta_{k+1}-\theta^{*}\|+\Big(3L_{x\theta}^{\Phi}D_{x}+\rho L_{f,\theta}JD_{f}(\theta^{*})\Big)\|\theta_{k}-\theta^{*}\|$
$\quad+L_{F,\theta}\big(\|\theta_{k}-\theta^{*}\|^{2}+\|\theta_{k-1}-\theta^{*}\|^{2}\big)+\frac{L_{\lambda\theta}^{\Phi}}{2\alpha_{k}}\|\theta_{k}-\theta^{*}\|^{2},$ (31)
where the last line uses $\|\theta_{k}-\theta_{k-1}\|^{2}\leq 2\|\theta_{k}-\theta^{*}\|^{2}+2\|\theta_{k-1}-\theta^{*}\|^{2}$. Summing over $k$ and reindexing the $\|\theta_{k-1}-\theta^{*}\|^{2}$ term, we obtain
$\sum_{k=0}^{\infty}R_{k}$ $\leq L_{F,\theta}D_{x}\sum_{k=1}^{\infty}\|\theta_{k}-\theta^{*}\|+\Big(3L_{x\theta}^{\Phi}D_{x}+\rho L_{f,\theta}JD_{f}(\theta^{*})\Big)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|$
$\quad+L_{F,\theta}\Big(\|\theta_{0}-\theta^{*}\|^{2}+2\sum_{k=1}^{\infty}\|\theta_{k}-\theta^{*}\|^{2}\Big)+\frac{L_{\lambda\theta}^{\Phi}}{2}\sum_{k=0}^{\infty}\frac{\|\theta_{k}-\theta^{*}\|^{2}}{\alpha_{k}}:=C_{R}.$
@@
Since $0<\alpha_{k}<1$, by the learning–error summability assumption in the statement of the lemma, the right-hand side of the above inequality is finite, i.e., $\sum_{k=0}^{\infty}R_{k}\leq C_{R}<\infty.$ Recall that, we can re-write (28) as follows
$\|\lambda_{k+1}-\lambda^{*}\|^{2}\leq(1+\alpha^{\prime}_{k})\|\lambda_{k}-\lambda^{*}\|^{2}+\beta_{k},$
where $\alpha^{\prime}_{k}=\frac{\rho L^{\Phi}_{\lambda\theta}\alpha_{k}}{1-\rho L^{\Phi}_{\lambda\theta}\alpha_{k}}$ and $\beta_{k}=\frac{2\rho}{1-\rho L^{\Phi}_{\lambda\theta}\alpha_{k}}\bigl(T_{k}+R_{k}\bigr)$. Fix any $\delta>0$ and $c\in(0,\frac{1}{2})$, selecting $\alpha_{k}=\frac{c}{(k+1)^{1+\delta}}$ and $\rho\leq\frac{1}{L^{\Phi}_{\lambda\theta}}$ implies that $t_{k}=:\rho L_{\lambda\theta}^{\Phi}\alpha_{k}\in(0,\tfrac{1}{2}]$, hence $\alpha_{k}^{\prime}=\frac{t_{k}}{1-t_{k}}\leq 2t_{k}$. Therefore, $\sum_{k=0}^{\infty}\alpha_{k}^{\prime}<\infty,$ and $\dfrac{2\rho}{1-\rho L_{\lambda\theta}^{\Phi}\alpha_{k}}\leq 4\rho$, which we conclude that $\sum_{k=0}^{\infty}\beta_{k}\leq 4\rho\sum_{k=0}^{\infty}\big(T_{k}+{R}_{k}\big)<\infty$. Finally, from [7, Lemma 5.31] we conclude that $\{\lambda_{k}\}$ converges to a finite limit, hence, it is uniformly bounded. ∎
Appendix D Proof of Theorem 4.4
@@
From (B), for every $k\geq 0$, $x\in X$, $\lambda\in\mathbb{R}^{J}_{+}$, we have
$(x_{k+1}-x)^{\top}F(x,\theta^{*})+\sum_{j=1}^{J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\Phi_{\rho}(x,\lambda_{k},\theta^{*})$ $\leq\mathcal{T}_{k}(x)+\mathcal{R}_{k}(x,\lambda)$
$\qquad+\frac{1}{2\rho}\bigl(\|\lambda_{k}-\lambda\|^{2}-\|\lambda_{k+1}-\lambda\|^{2}\bigr),$ (32)
@@
$\mathcal{T}_{k}(x)$ $=\frac{1}{2\gamma}\bigl(\|x_{k}-x\|^{2}-\|x_{k+1}-x\|^{2}\bigr)+\frac{L_{F,x}}{2}\big(\|x_{k}-x_{k-1}\|^{2}-\|x_{k+1}-x_{k}\|^{2}\big)$
@@
$\mathcal{R}_{k}(x,\lambda)$ $=\Big(\tfrac{L^{\Phi}+2L_{F,x}+L_{F,\theta}}{2}-\tfrac{1}{2\gamma}\Big)\|x_{k+1}-x_{k}\|^{2}+\tfrac{L_{F,\theta}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}$
$\quad+\tfrac{L_{F,\theta}}{2}\big(\|\theta_{k}-\theta_{k-1}\|^{2}-\|\theta_{k+1}-\theta_{k}\|^{2}\big)+L_{F,\theta}\|\theta_{k+1}-\theta^{*}\|\|x_{k+1}-x\|$
@@
$\quad+\rho L_{f,\theta}\|\theta_{k}-\theta^{*}\|\sum_{j\in J_{k}^{+}}\big|f_{j}(x_{k+1},\theta^{*})\big|.$
@@
Moving $\Phi_{\rho}(x,\lambda_{k},\theta^{*})$ to the right in (D), then summing over $k=0,\cdots,K-1$ and dividing by $K$, and using convexity of each $f_{j}(\cdot)$ to pass the average inside, we get
$(\bar{x}_{K}-x)^{\top}F(x,\theta^{*})+f(\bar{x}_{K},\theta^{*})^{\top}\lambda\leq\frac{1}{K}\sum_{k=0}^{K-1}\Phi_{\rho}(x,\lambda_{k},\theta^{*})+\frac{1}{K}\Delta_{K}(x,\lambda),$ (33)
where
$\Delta_{K}=\sum_{k=0}^{K-1}\bigl(\mathcal{T}_{k}(x)+\mathcal{R}_{k}(x,\lambda)\bigr)+\frac{1}{2\rho}\bigl(\|\lambda_{0}-\lambda\|^{2}-\|\lambda_{K}-\lambda\|^{2}\bigr).$ (34)
Then by the continuity on compact $X$ choose $D_{f}(\theta^{*})$ so that $\sup_{x\in X}|f_{j}(x,\theta^{*})|\leq D_{f}(\theta^{*})$. Then choosing the stepsize $\gamma\in\bigl(0,(L^{\Phi}+2L_{F,x}+L_{F,\theta})^{-1}\bigr)$ and $\rho_{k}=\rho$, the coefficient of $\|x_{k+1}-x_{k}\|^{2}$ in $\mathcal{T}_{k}$ is nonpositive. Telescoping in $x$, $\theta$, and the $r_{k}$-pair then yields a finite constant

### v3

@@ 5 Numerical Experiment
Using the above inequality in (D), multiplying both sides by $2\rho_{k}$ and rearranging the terms, we obtain
$\left(1-\rho_{k}L^{\Phi}_{\lambda\theta}\alpha_{k}\right)\|\lambda_{k+1}-\lambda^{*}\|^{2}$ $\leq\|\lambda_{k}-\lambda^{*}\|^{2}+\rho_{k}\left(\rho_{k}C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}-\frac{1}{\gamma_{k}}\right)\|x_{k+1}-x_{k}\|^{2}$
$\qquad+\rho_{k}C_{2}\|\lambda_{k}-\lambda^{*}\|\|x_{k+1}-x_{k}\|^{2}+2\rho_{k}\left(T_{k}+R_{k}\right),$ (31)
where
$T_{k}$ $\triangleq\tfrac{1}{2\gamma_{k}}\left(\|x_{k}-x^{*}\|^{2}-\|x_{k+1}-x^{*}\|^{2}\right)+\tfrac{L_{F,x}}{2}\left(\|x_{k}-x_{k-1}\|^{2}-\|x_{k+1}-x_{k}\|^{2}\right)+\langle r_{k+1},x_{k+1}-x^{*}\rangle-\langle r_{k},x_{k}-x^{*}\rangle,$
$R_{k}$ $\triangleq L_{F,\theta}\|\theta_{k+1}-\theta^{*}\|\|x_{k+1}-x^{*}\|+L^{\Phi}_{x\theta}\|\theta_{k}-\theta^{*}\|\big(\|x_{k+1}-x_{k}\|+\|x_{k}-x^{*}\|\big)+\tfrac{L_{F,\theta}}{2}\|\theta_{k}-\theta_{k-1}\|^{2}$
$\quad+\tfrac{L^{\Phi}_{\lambda\theta}}{2\alpha_{k}}\|\theta_{k}-\theta^{*}\|^{2}+\frac{J\rho_{k}}{2}L^{2}_{f,\theta}\|\theta_{k}-\theta^{*}\|^{2}+2L_{f,\theta}(J\rho_{k}D_{f}+\sqrt{J}\|\lambda_{k}\|)\|\theta_{k}-\theta^{*}\|.$
Assume now that $0<\rho_{k}L_{\lambda\theta}^{\Phi}\alpha_{k}<1$. Then, dividing both sides of (D) by $\left(1-\rho_{k}L^{\Phi}_{\lambda\theta}\alpha_{k}\right)$ we get
$\|\lambda_{k+1}-\lambda^{*}\|^{2}$ $\leq\frac{1}{\left(1-\rho_{k}L^{\Phi}_{\lambda\theta}\alpha_{k}\right)}\|\lambda_{k}-\lambda^{*}\|^{2}+\frac{\rho_{k}}{\left(1-\rho_{k}L^{\Phi}_{\lambda\theta}\alpha_{k}\right)}C_{2}\|\lambda_{k}-\lambda^{*}\|\|x_{k+1}-x_{k}\|^{2}$
$\qquad+\frac{\rho_{k}}{\left(1-\rho_{k}L^{\Phi}_{\lambda\theta}\alpha_{k}\right)}\left(\rho_{k}C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}-\frac{1}{\gamma_{k}}\right)\|x_{k+1}-x_{k}\|^{2}$
$\qquad+\frac{2\rho_{k}}{\left(1-\rho_{k}L^{\Phi}_{\lambda\theta}\alpha_{k}\right)}\left(T_{k}+R_{k}\right).$ (32)
We now provide upper bounds the auxiliary terms $T_{k}$ and $R_{k}$. For $T_{k}$ summing from $k=0,\dots,K-1$, using $x_{-1}=x_{0}$ and $r_{0}=0$, and $\gamma_{k}=\gamma$ we get
@@
From the definition of $T_{k}$ and telescoping with $x_{-1}=x_{0}$ and $r_{0}=0$ in (D), we have
@@
Choosing $\gamma\in\Big(0,(L_{F,x}+L_{F,\theta})^{-1}\Big)$ the coefficient of $\|x_{K}-x^{*}\|^{2}$ on the right-hand side of (34) is non-positive and may be discarded. Hence, $\sum_{k=0}^{K-1}T_{k}\leq\frac{1}{2\gamma}\|x_{0}-x^{*}\|^{2}+\frac{L_{F,\theta}}{2}\|\theta_{K}-\theta_{K-1}\|^{2}.$ By the compactness of $\Theta$, $\|\theta_{K}-\theta_{K-1}\|\leq\|\theta_{K}-\theta^{*}\|+\|\theta_{K-1}-\theta^{*}\|\leq 2D_{\Theta}$, it follows that
@@
where $C_{T}<\infty$ is independent of $K$. Therefore, for every $K\geq 1$, $\sum_{k=0}^{K-1}T_{k}\leq C_{T}.$
Next, we estimate $R_{k}$. Again, by the compactness of $X$, $\|x_{k+1}-x^{*}\|\leq D_{x}$ and $\|x_{k+1}-x_{k}\|\leq\|x_{k+1}-x^{*}\|+\|x_{k}-x^{*}\|\leq 2D_{x}$. Therefore, for every $k\geq 0$,
@@
$\quad+\frac{L_{\lambda\theta}^{\Phi}}{2\alpha_{k}}\|\theta_{k}-\theta^{*}\|^{2}+\frac{J\rho_{k}}{2}L_{f,\theta}^{2}\|\theta_{k}-\theta^{*}\|^{2}+2L_{f,\theta}\big(J\rho_{k}D_{f}+\sqrt{J}\|\lambda_{k}\|\big)\|\theta_{k}-\theta^{*}\|$
$\leq L_{F,\theta}D_{X}\|\theta_{k+1}-\theta^{*}\|+\Big(3L_{x\theta}^{\Phi}D_{X}+2J\rho_{k}D_{f}L_{f,\theta}\Big)\|\theta_{k}-\theta^{*}\|+\frac{L_{F,\theta}}{2}\|\theta_{k}-\theta_{k-1}\|^{2}$
@@
$\quad+\left(\frac{L_{\lambda\theta}^{\Phi}}{2\alpha_{k}}+\frac{J\rho_{k}}{2}L_{f,\theta}^{2}\right)\|\theta_{k}-\theta^{*}\|^{2}+2\sqrt{J}L_{f,\theta}\|\lambda_{k}\|\|\theta_{k}-\theta^{*}\|.$ (35)
Using $\|\theta_{k}-\theta_{k-1}\|^{2}\leq 2\|\theta_{k}-\theta^{*}\|^{2}+2\|\theta_{k-1}-\theta^{*}\|^{2},$ we further obtain
$R_{k}$ $\leq L_{F,\theta}D_{X}\|\theta_{k+1}-\theta^{*}\|+\Big(3L_{x\theta}^{\Phi}D_{X}+2J\rho_{k}D_{f}L_{f,\theta}\Big)\|\theta_{k}-\theta^{*}\|$
$\quad+L_{F,\theta}\big(\|\theta_{k}-\theta^{*}\|^{2}+\|\theta_{k-1}-\theta^{*}\|^{2}\big)+\left(\frac{L_{\lambda\theta}^{\Phi}}{2\alpha_{k}}+\frac{J\rho_{k}}{2}L_{f,\theta}^{2}\right)\|\theta_{k}-\theta^{*}\|^{2}$
$\quad+2\sqrt{J}L_{f,\theta}\|\lambda_{k}\|\,\|\theta_{k}-\theta^{*}\|.$ (36)
From the induction hypothesis, we have that $\|\lambda_{k}-\lambda^{*}\|\leq D_{\Lambda}$, $k=0,1,\dots,K-1$, hence, by the triangle inequality, $\|\lambda_{k}\|\leq\|\lambda_{k}-\lambda^{*}\|+\|\lambda^{*}\|\leq D_{\Lambda}+\|\lambda^{*}\|$, $k=0,1,\dots,K-1.$ Therefore, for $k=0,1,\dots,K-1$, (D) implies
$R_{k}$ $\leq L_{F,\theta}D_{X}\|\theta_{k+1}-\theta^{*}\|+\Big(3L_{x\theta}^{\Phi}D_{X}+2J\rho D_{f}L_{f,\theta}+2\sqrt{J}L_{f,\theta}(D_{\Lambda}+\|\lambda^{*}\|)\Big)\|\theta_{k}-\theta^{*}\|$
$\quad+L_{F,\theta}\big(\|\theta_{k}-\theta^{*}\|^{2}+\|\theta_{k-1}-\theta^{*}\|^{2}\big)+\left(\frac{L_{\lambda\theta}^{\Phi}}{2\alpha_{k}}+\frac{J\rho}{2}L_{f,\theta}^{2}\right)\|\theta_{k}-\theta^{*}\|^{2}.$ (37)
Summing from $k=0$ to (K1)(K-1) and noting that $\theta_{-1}=\theta_{0}$, we obtain
$\sum_{k=0}^{K-1}R_{k}$ $\leq L_{F,\theta}D_{X}\sum_{k=1}^{K}\|\theta_{k}-\theta^{*}\|+\Big(3L_{x\theta}^{\Phi}D_{X}+2J\rho D_{f}L_{f,\theta}+2\sqrt{J}L_{f,\theta}(D_{\Lambda}+\|\lambda^{*}\|)\Big)\sum_{k=0}^{K-1}\|\theta_{k}-\theta^{*}\|$
$\quad+L_{F,\theta}\Big(\|\theta_{0}-\theta^{*}\|^{2}+2\sum_{k=0}^{K-1}\|\theta_{k}-\theta^{*}\|^{2}\Big)+\sum_{k=0}^{K-1}\left(\frac{L_{\lambda\theta}^{\Phi}}{2\alpha_{k}}+\frac{J\rho}{2}L_{f,\theta}^{2}\right)\|\theta_{k}-\theta^{*}\|^{2}=:C_{R}(D_{\Lambda}).$
@@
Since $\alpha_{k}=\frac{1}{(k+2)^{2}}\in(0,1)$, by the learning–error summability assumption, the right-hand side is finite for each $K$, and thus $\sum_{k=0}^{K-1}R_{k}\leq C_{R}(D_{\Lambda}).$
Now define $\delta_{k}=\frac{1}{1-\rho_{k}L_{\lambda\theta}^{\Phi}\alpha_{k}}$, then (D), becomes
$\|\lambda_{k+1}-\lambda^{*}\|^{2}$ $\leq\delta_{k}\|\lambda_{k}-\lambda^{*}\|^{2}+\delta_{k}\rho_{k}C_{2}\|\lambda_{k}-\lambda^{*}\|\|x_{k+1}-x_{k}\|^{2}$
$\qquad+\delta_{k}\rho_{k}\left(\rho_{k}C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}-\frac{1}{\gamma}\right)\|x_{k+1}-x_{k}\|^{2}+2\delta_{k}\rho_{k}(T_{k}+R_{k}).$ (38)
Next, define the sequence $\{t_{k}\}_{k\geq 0}$ such that $t_{0}=1$ and $t_{k+1}=\frac{t_{k}}{\delta_{k}}\text{ for }k\geq 0$. Based on the parameter selections $\rho_{k}=\rho=\frac{1}{L_{\lambda\theta}^{\Phi}}$ and $\alpha_{k}=\frac{1}{(k+2)^{2}}$, we conclude that $t_{k}=\frac{t_{0}(k+2)}{2(k+1)}$ which implies that $t_{k}\in[\frac{1}{2},1]$ for any $k\geq 0$. Multiplying both sides of (D) by $t_{k+1}$, we obtain
$t_{k+1}\|\lambda_{k+1}-\lambda^{*}\|^{2}$ $\leq t_{k}\|\lambda_{k}-\lambda^{*}\|^{2}+\rho_{k}t_{k}C_{2}\|\lambda_{k}-\lambda^{*}\|\|x_{k+1}-x_{k}\|^{2}$
$\qquad+\rho_{k}t_{k}\left(\rho_{k}C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}-\frac{1}{\gamma}\right)\|x_{k+1}-x_{k}\|^{2}+2\rho_{k}t_{k}(T_{k}+R_{k}).$ (39)
We prove the boundedness of the dual iterates by induction, showing that there exists $D_{\Lambda}>0$ such that $\|\lambda_{k}-\lambda^{*}\|\leq D_{\Lambda}$ for all $k\geq 0.$ Assume, as the induction hypothesis, that $\|\lambda_{k}-\lambda^{*}\|\leq D_{\Lambda}\quad\text{for all }k=0,1,\dots,K-1.$ Hence, we get from (D)
$t_{k+1}\|\lambda_{k+1}-\lambda^{*}\|^{2}$ $\leq t_{k}\|\lambda_{k}-\lambda^{*}\|^{2}+\rho_{k}t_{k}\left(\rho_{k}C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}+C_{2}D_{\Lambda}-\frac{1}{\gamma}\right)\|x_{k+1}-x_{k}\|^{2}$
$\quad+2\rho_{k}t_{k}(T_{k}+R_{k}).$ (40)
Now, for any $\rho_{k}\equiv\rho>0$, let us choose $\gamma$ such that $\gamma\leq(\rho C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}+C_{2}D_{\Lambda})^{-1}$, then the coefficient of $\|x_{k+1}-x_{k}\|^{2}$ in (D) is non-positive and may be discarded. Therefore, we have
$t_{k+1}\|\lambda_{k+1}-\lambda^{*}\|^{2}\leq t_{k}\|\lambda_{k}-\lambda^{*}\|^{2}+2\rho t_{k}(T_{k}+R_{k}).$
Summing this inequality from $k=0$ to K1K-1 yields
$t_{K}\|\lambda_{K}-\lambda^{*}\|^{2}\leq t_{0}\|\lambda_{0}-\lambda^{*}\|^{2}+2\rho\sum_{k=0}^{K-1}t_{k}(T_{k}+R_{k}).$
Since $t_{k}\in[\frac{1}{2},1]$ for all $k\geq 0$, it follows that
$\frac{1}{2}\|\lambda_{K}-\lambda^{*}\|^{2}$ $\leq\|\lambda_{0}-\lambda^{*}\|^{2}+2\rho\sum_{k=0}^{K-1}(T_{k}+R_{k})$
$\leq\|\lambda_{0}-\lambda^{*}\|^{2}+2\rho\bigl(C_{T}+C_{R}(D_{\Lambda})\bigr),$
where the last inequality follows from the fact that $\sum_{k=0}^{K-1}T_{k}\leq C_{T}$ and $\sum_{k=0}^{K-1}R_{k}\leq C_{R}(D_{\Lambda}).$ Thus, we have
$\|\lambda_{K}-\lambda^{*}\|^{2}\leq 2\|\lambda_{0}-\lambda^{*}\|^{2}+4\rho\left(C_{T}+C_{R}(D_{\Lambda})\right)=2\|\lambda_{0}-\lambda^{*}\|^{2}+\frac{2\rho}{\gamma}\|x_{0}-x^{*}\|^{2}+8\rho L_{F,\theta}D_{\Theta}^{2}+4\rho C_{R}\left(D_{\Lambda}\right),$
where in the last equality we used the definition of $C_{T}$. Since both $\gamma=\mathcal{O}(D_{\Lambda})$ and $C_{R}\left(D_{\Lambda}\right)=\mathcal{O}(D_{\Lambda})$, selecting $D_{\Lambda}\triangleq B+\sqrt{A}$, where $A\triangleq 2\|\lambda_{0}-\lambda^{*}\|^{2}+2\rho\|x_{0}-x^{*}\|^{2}\bigl(\rho C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}\bigr)+8\rho L_{F,\theta}D_{\Theta}^{2}+4\rho\Bigg(\Bigl(L_{F,\theta}D_{X}+3L_{x\theta}^{\Phi}D_{X}+2J\rho D_{f}L_{f,\theta}+2\sqrt{J}L_{f,\theta}\|\lambda^{*}\|\Bigr)\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|+L_{F,\theta}\Bigl(D_{\Theta}^{2}+2\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|^{2}\Bigr)+\frac{L_{\lambda\theta}^{\Phi}}{2}\sum_{k=0}^{\infty}\frac{\|\theta_{k}-\theta^{*}\|^{2}}{\alpha_{k}}+\frac{J\rho}{2}L_{f,\theta}^{2}\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|^{2}\Bigg)$ and $B\triangleq 2\rho C_{2}\|x_{0}-x^{*}\|^{2}+8\rho\sqrt{J}\,L_{f,\theta}\sum_{k=0}^{\infty}\|\theta_{k}-\theta^{*}\|$, implies that $\|\lambda_{K}-\lambda^{*}\|^{2}\leq D_{\Lambda}^{2}$ which completes the induction, hence, the result is proved. ∎
Appendix E Proof of Theorem 4.5
@@
Proof From (C), for every $k\geq 0$, $x\in X$, $\lambda\in\mathbb{R}^{J}_{+}$, we have
$(x_{k+1}-x)^{\top}F(x,\theta^{*})+\sum_{j=1}^{J}\lambda^{(j)}f_{j}(x_{k+1},\theta^{*})-\Phi_{\rho}(x,\lambda_{k},\theta^{*})\leq\mathcal{T}_{k}(x)+\mathcal{R}_{k}(x,\lambda)+\frac{1}{2\rho}\bigl(\|\lambda_{k}-\lambda\|^{2}-\|\lambda_{k+1}-\lambda\|^{2}\bigr),$ (41)
@@
$\mathcal{T}_{k}(x)$ $\triangleq\frac{1}{2\gamma}\bigl(\|x_{k}-x\|^{2}-\|x_{k+1}-x\|^{2}\bigr)+\frac{L_{F,x}}{2}\big(\|x_{k}-x_{k-1}\|^{2}-\|x_{k+1}-x_{k}\|^{2}\big)$
@@
$\mathcal{R}_{k}(x,\lambda)$ $\triangleq\left(\frac{\rho_{k}C_{1}+\|\lambda^{*}\|C_{2}+2L_{F,x}+L_{F,\theta}}{2}-\frac{1}{2\gamma_{k}}\right)\|x_{k+1}-x_{k}\|^{2}+\frac{C_{2}}{2}\|\lambda_{k}-\lambda^{*}\|\|x_{k+1}-x_{k}\|^{2}$
$\quad+\tfrac{L_{F,\theta}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}+\tfrac{L_{F,\theta}}{2}\big(\|\theta_{k}-\theta_{k-1}\|^{2}-\|\theta_{k+1}-\theta_{k}\|^{2}\big)+L_{F,\theta}\|\theta_{k+1}-\theta^{*}\|\|x_{k+1}-x\|$
@@
$\quad+\frac{J\rho}{2}L_{f,\theta}^{2}\|\theta_{k}-\theta^{*}\|^{2}+2L_{f,\theta}(J\rho D_{f}+\sqrt{J}\|\lambda_{k}\|)\|\theta_{k}-\theta^{*}\|.$
@@
Moving $\Phi_{\rho}(x,\lambda_{k},\theta^{*})$ to the right in (41), then summing over $k=0,\cdots,K-1$ and dividing by $K$, and using convexity of each $f_{j}(\cdot,\theta^{*})$ to pass the average inside, we get
$(\bar{x}_{K}-x)^{\top}F(x,\theta^{*})+f(\bar{x}_{K},\theta^{*})^{\top}\lambda$ $\leq\frac{1}{K}\sum_{k=0}^{K-1}\Phi_{\rho}(x,\lambda_{k},\theta^{*})+\frac{1}{K}\sum_{k=0}^{K-1}\bigl(\mathcal{T}_{k}(x)+\mathcal{R}_{k}(x,\lambda)\bigr)$
$\quad+\frac{1}{2\rho}\bigl(\|\lambda_{0}-\lambda\|^{2}-\|\lambda_{K}-\lambda\|^{2}\bigr).$ (42)
Considering $\sum_{k=0}^{K-1}\mathcal{T}_{k}(x)$, telescoping in the $x$-terms and the $r_{k}$-pair, we have
$\sum_{k=0}^{K-1}\mathcal{T}_{k}(x)=\frac{1}{2\gamma}\left(\|x_{0}-x\|^{2}-\|x_{K}-x\|^{2}\right)-\frac{L_{F,x}}{2}\|x_{K}-x_{K-1}\|^{2}+\big\langle r_{K},x_{K}-x\big\rangle.$
Note that $\|r_{K}\|\leq L_{F,x}\|x_{K}-x_{K-1}\|+L_{F,\theta}\|\theta_{K}-\theta_{K-1}\|$, therefore $\big\langle r_{K},x_{K}-x\big\rangle\leq\frac{L_{F,x}}{2}\|x_{K}-x_{K-1}\|^{2}+\frac{L_{F,x}}{2}\|x_{K}-x\|^{2}+L_{F,\theta}\|\theta_{K}-\theta_{K-1}\|\|x_{K}-x\|$. Substituting this into the previous equality, we have
$\sum_{k=0}^{K-1}\mathcal{T}_{k}(x)=\frac{1}{2\gamma}\|x_{0}-x\|^{2}+\left(\frac{L_{F,x}}{2}-\frac{1}{2\gamma}\right)\|x_{K}-x\|^{2}+L_{F,\theta}\|\theta_{K}-\theta_{K-1}\|\|x_{K}-x\|.$
Choosing $\gamma\in\left(0,L_{F,x}^{-1}\right)$, the coefficient of $\|x_{K}-x\|^{2}$ is non-positive and can be discarded. Since $x,x_{o},x_{K}\in X$ and $\theta_{K},\theta_{K-1}\in\Theta$, we have $\|x_{0}-x\|\leq D_{X},\quad\|x_{K}-x\|\leq D_{X}$, and $\|\theta_{K}-\theta_{K-1}\|\leq 2D_{\Theta}$. Therefore,